# Adaptive Central Frequencies Locally Competitive Algorithm for Speech

### v1

@@ II-A Locally Competitive Algorithm (LCA)
where $\tau$ is the time constant of each neuron, $\boldsymbol{p}$ is the input signal projection on the dictionary, i.e., $\boldsymbol{p}=\boldsymbol{D}^{T}\boldsymbol{s}$, $\boldsymbol{v}$ is the membrane potential vector, and $\boldsymbol{I}$ is the identity matrix. Mainly, the evolution of $\boldsymbol{v}$ over time depends on the input intensity $\boldsymbol{p}$ as an integration term and on $-\boldsymbol{v}$ as a leak term which makes these neurons behave like leaky integrators. Membrane potentials exceeding the threshold $\lambda$ –––the same as the Lagrange multiplier in (2)––– produce activations corresponding to the coefficients $\boldsymbol{a}$. Each activated neuron inhibits all others through horizontal connections, whose weights are calculated as $\boldsymbol{D}^{T}\boldsymbol{D}-\boldsymbol{I}$. For each neuron $m$, the activation $a_{m}$ is a non-linearity $T_{\lambda}$ that can be sigmoidal or –––as we used in this work––– the hard thresholding function applied to the potential $v_{m}$ of neuron $m$:
@@ II-C Adaptive Locally Competitive Algorithm (ALCA)
where |||| is the ”OR” operator and $\boldsymbol{c}$, $\boldsymbol{b}$, and $\boldsymbol{l}$ are vectors containing the parameters of the filters $c_{i}$, $b_{i}$, and $l_{i}$, respectively. A trade-off parameter $\alpha$ is introduced to decouple LCA’s neurons threshold $\lambda$ in (4) from the Lagrange multiplier $\lambda$ in (2). This relationship enables the implementation of the gradient descent, allowing the gradient flow from the membrane potential to backpropagate further towards the Gammachirp’s parameters [10].
@@ II-D Adaptation of Central Frequency Distribution with ALCA (ALCA-CF)
$\frac{\partial E}{\partial\boldsymbol{f}}=\frac{\partial\frac{1}{2}||\boldsymbol{D}\boldsymbol{a}-\boldsymbol{s}||^{2}}{\partial\boldsymbol{D}\boldsymbol{a}}\frac{\partial\boldsymbol{D}\boldsymbol{a}}{\partial\boldsymbol{D}}\frac{\partial\boldsymbol{D}}{\partial\boldsymbol{f}}+\alpha\lambda\frac{dS(\boldsymbol{a})}{d\boldsymbol{a}}\frac{d\boldsymbol{a}}{d\boldsymbol{v}}\frac{\partial\boldsymbol{v}}{\partial\boldsymbol{D}}\frac{\partial\boldsymbol{D}}{\partial\boldsymbol{f}}$
@@ II-E Datasets and Representations
[…]
@@
For ALCA, the adaptation hyper-parameters were optimized in [8]. In this article, we use the selected values from [8]. We perform the same hyper-parameter optimization for ALCA-CF but we characterize the central frequencies with their own backpropagation learning rate lr-cf since their range is from 0 to the Nyquist frequency which is significantly different from the ranges of modulation parameters. The search interval of lr-cf is $[10^{-6},10^{2}]$ on a logarithmic scale.

### v2

@@ II-A Locally Competitive Algorithm (LCA)
where $\tau$ is the time constant of each neuron, $\boldsymbol{p}$ is the input signal projection on the dictionary, i.e., $\boldsymbol{p}=\boldsymbol{D}^{T}\boldsymbol{s}$, $\boldsymbol{v}$ is the membrane potential vector, and $\boldsymbol{I}$ is the identity matrix. Mainly, the evolution of $\boldsymbol{v}$ over time depends on the input intensity $\boldsymbol{p}$ as an integration term and on $\boldsymbol{v}$ as a leak term which makes these neurons behave like leaky integrators. Membrane potentials exceeding the threshold $\lambda$ –––the same as the Lagrange multiplier in (2)––– produce activations corresponding to the coefficients $\boldsymbol{a}$. Each activated neuron inhibits all others through horizontal connections, whose weights are calculated as $\boldsymbol{D}^{T}\boldsymbol{D}-\boldsymbol{I}$. For each neuron $m$, the activation $a_{m}$ is a non-linearity $T_{\lambda}$ that can be sigmoidal or –––as we used in this work––– the hard thresholding function applied to the potential $v_{m}$ of neuron $m$:
@@ II-C Adaptive Locally Competitive Algorithm (ALCA)
where $|$ is the ”OR” operator and $\boldsymbol{c}$, $\boldsymbol{b}$, and $\boldsymbol{l}$ are vectors containing the parameters of the filters $c_{i}$, $b_{i}$, and $l_{i}$, respectively. A trade-off parameter $\alpha$ is introduced to decouple LCA’s neurons threshold $\lambda$ in (4) from the Lagrange multiplier $\lambda$ in (2). This relationship enables the implementation of the gradient descent, allowing the gradient flow from the membrane potential to backpropagate further towards the Gammachirp’s parameters [10].
@@ II-D Adaptation of Central Frequency Distribution with ALCA (ALCA-CF)
[…]
@@ II-E Datasets and Representations
This study focuses on speech processing using two datasets. The first is Heidelberg Digits (HD) [9], which contains approximately $10\thinspace{}420$ recordings of spoken digits (099) in English and German from $12$ speakers, sampled at $48$ kHz. The dataset is divided into training and test sets. Two speakers exclusively reserved for testing, and $5\%$ of test samples drawn from speakers in the training set to evaluate speaker generalization. The second is Google Speech Commands (SC) [19], consisting of $105\thinspace{}829$ one-second audio files sampled at $16$ kHz, representing $35$ classes of English words spoken by $1\thinspace{}864$ speakers. We use version $2$ of this dataset, partitioned into training, testing, and validation sets using the hashing function from [9].
@@
For ALCA, the adaptation hyper-parameters were optimized in [8]. In this article, we use the selected values from [8]. We perform the same hyper-parameter optimization for ALCA-CF but we characterize the central frequencies with their own backpropagation learning rate lr-cf since their range is from 0 to the Nyquist frequency which is significantly different from the ranges of modulation parameters. The search interval of lr-cf is $10^{-6},10^{2}]$ on a logarithmic scale.